OpenStudy (anonymous):

-38 e^x -2 = -39 + 18 e^x Find the exact value of x. No decimal entries allowed.

OpenStudy (xishem):

Attempts?

OpenStudy (anonymous):

No solution mate

OpenStudy (anonymous):

yeah, I just want to know where to start do I combine?

OpenStudy (anonymous):

i'm kind of lost on this one.

OpenStudy (xishem):

Just to clarify, it's this, right? \[-38e^x-2=-39+18e^x\]

OpenStudy (anonymous):

yes it is

OpenStudy (xishem):

Then start by combining like terms.

OpenStudy (anonymous):

e^x= 37/56

OpenStudy (xishem):

Yep. And keep going.

OpenStudy (anonymous):

so now ln both sides?

OpenStudy (xishem):

Try it!

OpenStudy (anonymous):

Nice, thanks so whenever there is an e we just take the natural log of both sides to get rid of it? which property of logarithms is that?

OpenStudy (anonymous):

I got a question lets say for example we were to have 2^x=41 would we do the ln41/ln2 =x ?

OpenStudy (xishem):

\[y=\log_bx \\ b^y=x\]Logarithms and exponents are inverses of each other, so when the two are nested, they cancel.

OpenStudy (anonymous):

oh okay, that makes sense I gotta remember the properties exam tomorrow and final next Tuesday.

OpenStudy (xishem):

\[2^x=41\]In this case, you need to take the log(base 2) of both sides:\[\log_2(2^x)=\log_2(41)\]\[x=\log_2(41)=\frac{\log(41)}{\log(2)}\]

OpenStudy (anonymous):

why log base 2?

OpenStudy (xishem):

Because log_2 is the inverse of 2^.

OpenStudy (xishem):

It's just like when you have: \[2x=1\]To isolate x, you perform the INVERSE of the multiplication by 2 (division by 2):\[\frac{2x}{2}=\frac{1}{2}\]It's the same idea, just with logs and powers.

OpenStudy (anonymous):

That makes more sense now thanks.